An unbalanced four wire star connected load has a balanced voltage of 400v. The load are Z1= 4+j8, Z2= 3+j4, Z3= 15+j20 ohms.


A.) Find the line current

B.) Find the neutral current


1
Expert's answer
2021-01-18T01:13:46-0500

a) "I_L=\frac{U}{Z_1+Z_2+Z_3}=\frac{400}{4+3+15+j8+j4+j20}=5.84-j8.49" A

b)"I_L=\sqrt{3}\times I_n; I_n=\frac{I_L}{\sqrt{3}}=\frac{5.84-j8.49}{\sqrt{3}}=3.37-j4,9" A